# Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise

### v3

@@ III Proposed Joint CFO, PN and UE Position Estimation Algorithm
In previous works [27, 30], localization was based on frequency-domain signals, and different subcarriers’ signals were often assumed to be independent. However, the CFO and PN would result in ICI, which increases the difficulty in the estimation of the CFO, PN and the position of the UE. The estimation can only be conducted in the time domain. In this section, we propose an AO-based iterative algorithm, which employs the MAP criterion and GD algorithm to jointly estimate the normalized CFO, PN, and the UE position.
Define the parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, wherein $\bm{\theta}$ is a random variable and the others are deterministic. The posterior distribution of $\bm{\lambda}$, given the received signal $\bm{y}$, can be expressed as
@@
$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}),$ (8)
where $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, which is obtained by calculating the negative logarithmic function of (7) and then suppressing constants.
@@ III-A1 Approximation of PN vector
From the definition of $\bm{\lambda}$, many variables need to be jointly estimated at the UE. To lower the computational complexity of the estimation, we exploit the redundancy of the covariance matrix of the PN vector to approximate it and reduce variables requiring estimation.
@@
where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN variables, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transition matrix. We set $L=\frac{N}{2}$ in this paper. To derive $\bm{\Pi}$, we first perform eigenvalue decomposition (EVD) on $\bm{\Psi}$, i.e., $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$ and $\bm{e}$ collects the eigenvalues of $\bm{\Psi}$ arranged in descending order. Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times\sqrt{\bm{e}}$. In the following analysis, $\bm{\eta}$ is estimated, instead of $\bm{\theta}$.
@@ III-A2 Estimation of approximate PN vector
Set the partial derivative of (III-A2) regarding $\bm{\eta}$ to zero, i.e.,
@@ III-B CFO Estimation
To enhance the tractability of (14) and derive a closed-form solution, the first-order Taylor approximation is employed. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
$\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n}{N}\phi}\approx e^{j2\pi\frac{n}{N}\hat{\phi}}+\left(\phi-\hat{\phi}\right)\frac{j2\pi}{N}e^{j2\pi\frac{n}{N}\hat{\phi}}.$ (15)
@@
where $\left[\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]_{n,n}=\frac{j2\pi n}{N}e^{\frac{j2\pi n\hat{\phi}^{[q]}}{N}}$.
@@ III-C Position Estimation
For position estimation, we employ the GD algorithm. Due to the same structures in the respective derivatives of the objective function with regard to the three position parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them. The partial derivative of the objective function with regard to $\xi$ is given by
@@
$\frac{\partial\bm{h}}{\partial\xi}[n]$ $=\frac{\partial\bm{h}[n]}{\partial\xi}=-\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]$
×$\times e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}\left(\frac{1}{d_{ru}}+j2\pi\frac{nB}{Nc}\right)\frac{\partial d_{ru}}{\partial\xi},$ (20)
with $\frac{\partial d_{ru}}{\partial\xi}$, $\xi=d_{Ru},\phi_{az},\phi_{el}$, being
$\frac{\partial d_{ru}}{\partial d_{Ru}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})\sin\phi_{el}\cos\phi_{az}\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})\sin\phi_{el}\sin\phi_{az}\right.$
$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})\cos\phi_{el}\right],$ (21)
$\frac{\partial d_{ru}}{\partial\phi_{az}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})(-\sin\phi_{az}d_{Ru}\sin\phi_{el})\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{el}\cos\phi_{az}\right],$ (22)
@@
$\frac{\partial d_{ru}}{\partial\phi_{el}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})d_{Ru}\cos\phi_{az}\cos\phi_{el}\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{az}\cos\phi_{el}\right.$
$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
@@ III-D Overall Algorithm
The proposed joint estimation of the CFO, PN and UE position is summarized in Algorithm 1. Specifically, we initialize the parameters, including the initial estimation of the normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, the step size, $k_{p}$, the convergence accuracy for outer and inner loops, $\epsilon_{o}$ and $\epsilon_{i}$, and the maximum iteration number for the outer and inner loops, $N_{o}$ and $N_{i}$. After transforming the coordinates into the polar coordinates, we estimate the PN, CFO, and UE position alternately. Closed-form expressions are utilized to update the PN and CFO estimates in each iteration, while the GD algorithm is employed to estimate the UE position. This repeats until either the objective function $\mathcal{L}$ converges or the maximum number of iterations is reached. It is noteworthy that this AO-based algorithm takes the RIS phase shifts, i.e., $\bm{w}$, as input, which can affect the algorithm performance significantly. The configuration of the RIS phase shifts, $\bm{w}$, will be discussed in Section V.
@@
The computational complexity for updating $\bm{\eta}$, $\bm{\theta}$, $\phi$, and position parameters $d_{Rt}$, $\phi_{az}$, and $\phi_{el}$ in Algorithm 1 are $\mathcal{O}(L^{3}+N^{2}+NL^{2})$, $\mathcal{O}(NL)$, $\mathcal{O}(N^{2})$, and $\mathcal{O}(\max\left\{N^{3},N_{R}\right\})$, respectively. Suppose the inner GD algorithm for updating the UE position converges after $K_{i}$ iterations. Consequently, the computational complexity of Algorithm 1 is $\mathcal{O}(\max\left\{L^{3}+N^{2}+NL^{2},K_{i}N^{3},K_{i}N_{R}\right\})$ per iteration.
@@
2:  Initialization: $\hat{\phi}$, $\hat{\bm{p}}_{u}$, $k_{p}$, $\epsilon_{o}$, $\epsilon_{i}$, $N_{o}$, $N_{i}$;
@@
5:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\bm{\eta}}$ with $\hat{\phi}$ using (13);
@@
7:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\phi}$ with $\hat{\bm{\theta}}$ using (18);
@@
9:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
10:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
11:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
12:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration or $p\textgreater N_{i}$.
13:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration or $q\textgreater N_{o}$.
@@
15:  Output: $\hat{\phi}$, $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$.

### v4

@@ III Proposed Joint CFO, PN and UE Position Estimation Algorithm
In previous works [26, 29], localization was based on frequency-domain signals, and different subcarriers’ signals were often assumed to be independent. However, the CFO and PN can result in ICI, which increases the difficulty in the estimation of the CFO, PN, and UE position. The estimation can only be conducted in the time domain. In this section, we propose an AO-based iterative algorithm, which employs the MAP criterion and GD algorithm to jointly estimate the normalized CFO, PN, and the UE position. To address the new challenge of estimating the CFO and PN, the PN vector is approximated by exploiting the redundancy of its covariance matrix, and the posterior functions concerning the CFO and PN are linearized. Closed-form estimates are derived for both CFO and PN.
The proposed localization algorithm can serve as a prerequisite for tracking. Since the Doppler shift can be interpreted as part of CFO, by estimating the CFO, we can provide an estimate of the UE’s speed and improve tracking. For example, the UE position and CFO estimates provided by our proposed localization algorithm can be used as initialization inputs for existing tracking algorithms, e.g., [29].
Define the parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},{\color[rgb]{0,0,0}d_{\text{R}}},{\color[rgb]{0,0,0}\phi_{\text{az}}},{\color[rgb]{0,0,0}\phi_{\text{el}}}\right]^{\top}$, wherein $\bm{\theta}$ is a random variable and the others are deterministic. The $k$-th variable in $\bm{\lambda}$ is denoted as $\lambda_{k}$. The posterior distribution of $\bm{\lambda}$, given the received signal $\bm{y}$, can be expressed as
@@
$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\>-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta}),$ (9)
which is obtained by calculating the negative logarithmic function of (8) and then suppressing constants.
@@ III-A1 Approximation of PN vector
From the definition of $\bm{\lambda}$, many variables need to be jointly estimated at the UE. To lower the computational complexity of the estimation, we exploit the redundancy of the covariance matrix of the PN vector to approximate it and reduce variables requiring estimation. This approximation is based on the commonly used truncated singular value decomposition (SVD) method [36, 37].
@@
where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN variables, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transition matrix. To derive $\bm{\Pi}$, we first perform eigenvalue decomposition (EVD) on $\bm{\Psi}$. For the real symmetric matrix $\mathbf{\Psi}$, the EVD is equivalent to its SVD. Based on the EVD, we have $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$ and $\bm{e}$ collects the eigenvalues of $\bm{\Psi}$ arranged in descending order. Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times{\rm Diag}\left\{\tilde{\bm{e}}\right\}$, with $\tilde{\bm{e}}\triangleq\left[\sqrt{\bm{e}[0]},\sqrt{\bm{e}[1]},\cdots,\sqrt{\bm{e}[L-1]}\right]^{\top}$.
Define $\tilde{\mathbf{\Psi}}=\mathbf{\Pi}\mathbf{\Pi}^{\top}$. Thus, $\mathbf{\Pi}\bm{\eta}\sim\mathcal{N}(\bm{0}_{N\times 1},\tilde{\mathbf{\Psi}})$. $\tilde{\mathbf{\Psi}}$ can be rewritten as
$\tilde{\mathbf{\Psi}}$ $=\left[\mathbf{U}\right]_{:,0:L-1}{\rm Diag}\left\{\tilde{\bm{e}}\right\}\left\{\tilde{\bm{e}}\right\}^{\top}\left[\mathbf{U}\right]_{:,0:L-1}^{\top}$
$=\sum_{l=0}^{L-1}\bm{e}[l]\left[\mathbf{U}\right]_{:,l}\left[\mathbf{U}\right]_{:,l}^{\top},$ (11)
and is expected to approach $\mathbf{\Psi}$ when $L$ is sufficiently large since most of the eigenvalues of $\mathbf{\Psi}$ are close to zero. In this paper, we set $L=\frac{N}{2}$. In the following analysis, $\bm{\eta}$ is estimated, instead of $\bm{\theta}$.
@@ III-A2 Estimation of approximate PN vector
Setting the partial derivative of (III-A2) regarding $\bm{\eta}$ to zero,
@@ III-B CFO Estimation
To enhance the tractability of (16) and derive a closed-form solution, the first-order Taylor approximation is employed [38, 39]. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
$\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n}{N}\phi}\approx e^{j2\pi\frac{n}{N}\hat{\phi}}+\left(\phi-\hat{\phi}\right)\frac{j2\pi{\color[rgb]{0,0,0}n}}{N}e^{j2\pi\frac{n}{N}\hat{\phi}}.$ (17)
@@
where $\left[\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]_{n,n}=\frac{j2\pi n}{N}e^{\frac{j2\pi n\hat{\phi}^{[q]}}{N}}$. By defining $x=\phi-\hat{\phi}$, the approximation in (17) can be rewritten as
$f(x)=e^{j2\pi\frac{n}{N}x}\approx 1+\frac{j2\pi{\color[rgb]{0,0,0}n}}{N}x,$ (19)
which is essentially the first-order Taylor series expansion of $f(x)$ at $x=0$, with $x$ being the estimation accuracy of $\phi$. The Lagrange remainder of this Taylor expansion, which reflects the approximation accuracy, can be calculated as
${\rm R}_{1}(x)$ $=\frac{f^{{}^{\prime\prime}}(\xi)}{2}x^{2}=-2\pi^{2}\frac{n^{2}}{N^{2}}e^{j2\pi\frac{n}{N}\xi}x^{2},$ (20)
where $\xi\in(0,x)$. Since $|e^{j2\pi\frac{n}{N}\xi}|=1$, $|{\rm R}_{1}(x)|\leq 2\pi^{2}\frac{n^{2}}{N^{2}}x^{2}$. The remainder of the first-order Taylor expansion of $f(x)$ is upper bounded, and the upper bound depends on the estimation accuracy of $\phi$. As the estimation accuracy of $\phi$ improves, the Taylor approximation in (17) becomes increasingly accurate.
@@ III-C Position Estimation
The GD algorithm is a general method for searching maxima or minima in unconstrained optimization problems and has been widely used to solve localization problems in existing studies [40, 41, 42]. Under the given CFO and PN estimates, the MAP-based problem in (9) is an unconstrained optimization problem. We employ the GD algorithm to estimate UE position.
Due to the same structures in the respective derivatives of the objective function with regard to the three position parameters ${\color[rgb]{0,0,0}d_{\text{R}}}$, ${\color[rgb]{0,0,0}\phi_{\text{az}}}$, and ${\color[rgb]{0,0,0}\phi_{\text{el}}}$, we use the symbol $\xi$ to represent any one of them. The partial derivative of the objective function with regard to $\xi$ is given by
@@
$\frac{\partial\bm{h}}{\partial\xi}[n]$ $=\frac{\partial\bm{h}[n]}{\partial\xi}=-\sum_{r=0}^{{\color[rgb]{0,0,0}N_{\text{R}}-1}}{\color[rgb]{0,0,0}\sqrt{\rho_{{\text{aR}}}^{n,r}}}{\color[rgb]{0,0,0}\sqrt{\rho_{{\text{Ru}}}^{n,r}}}\bm{w}[r]$
×$\times e^{-j2\pi\frac{nB}{N}({\tau_{{\color[rgb]{0,0,0}\text{a}}r}}+{\color[rgb]{0,0,0}\tau_{r\text{u}}})}\left(\frac{1}{{\color[rgb]{0,0,0}d_{r\text{u}}}}+j2\pi\frac{nB}{Nc}\right)\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial\xi},$ (24)
with $\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial\xi}$, $\xi={\color[rgb]{0,0,0}d_{\text{R}}},{\color[rgb]{0,0,0}\phi_{\text{az}}},{\color[rgb]{0,0,0}\phi_{\text{el}}}$, being
$\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\frac{1}{{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}d_{r\text{u}}}}}\left[({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}-x_{r})\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\right.$
$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}-y_{r})\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}\right.$
$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}-z_{r})\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\right],$ (25)
$\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\frac{1}{{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}d_{r\text{u}}}}}\left[({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}-x_{r})(-\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}})\right.$
$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}-y_{r}){\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\right],$ (26)
@@
$\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\frac{1}{{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}d_{r\text{u}}}}}\left[({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}-x_{r}){\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\right.$
$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}-y_{r}){\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\right.$
$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}-z_{r})(-{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}})\right].$ (27)
@@ III-D Overall Algorithm
The proposed joint estimation of the CFO, PN and UE position is summarized in Algorithm 1. Specifically, we initialize the parameters, including the initial estimation of the normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, the step size $k_{\color[rgb]{0,0,0}\text{p}}$, the convergence accuracy requirements for outer and inner loops, $\epsilon_{\color[rgb]{0,0,0}\text{O}}$ and $\epsilon_{\color[rgb]{0,0,0}\text{I}}$, and the maximum numbers of iterations for the outer and inner loops, $N_{\color[rgb]{0,0,0}\text{O}}$ and $N_{\color[rgb]{0,0,0}\text{I}}$. Suppose that the range of the normalized CFO $\phi$ is known at the UE, and the initial $\hat{\phi}$ is chosen randomly within this range. The initial estimation of the UE position, $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, is obtained using the state of the art proposed in [26]. The other initial parameters are hyperparameters, and can be fine-tuned to balance both convergence speed and convergence accuracy.
Then, after transforming the coordinates into the polar coordinates, we estimate the PN, CFO, and UE position alternately. Closed-form expressions are utilized to update the PN and CFO estimates in each iteration, while the GD algorithm is employed to estimate the UE position. This repeats until either the objective function $\mathcal{L}$ converges or the maximum number of iterations is reached. Notably, this AO-based algorithm takes the RIS phase shifts, i.e., $\bm{w}$, as input, which can affect the algorithm performance significantly. The configuration of the RIS phase shifts, $\bm{w}$, will be discussed in Section V.
@@
The computational complexity for updating $\bm{\eta}$, $\bm{\theta}$, $\phi$, and position parameters ${\color[rgb]{0,0,0}d_{\text{R}}}$, ${\color[rgb]{0,0,0}\phi_{\text{az}}}$, and ${\color[rgb]{0,0,0}\phi_{\text{el}}}$ in Algorithm 1 are $\mathcal{O}(L^{3}+N^{2}+NL^{2})$, $\mathcal{O}(NL)$, $\mathcal{O}(N^{2})$, and $\mathcal{O}(\max\left\{N^{3},{\color[rgb]{0,0,0}N_{\text{R}}}\right\})$, respectively. Suppose the inner GD algorithm for updating the UE position converges after $K_{\color[rgb]{0,0,0}\text{I}}$ iterations. Consequently, the computational complexity of Algorithm 1 is $\mathcal{O}(\max\left\{L^{3}+N^{2}+NL^{2},K_{\color[rgb]{0,0,0}\text{I}}N^{3},K_{\color[rgb]{0,0,0}\text{I}}{\color[rgb]{0,0,0}N_{\text{R}}}\right\})$ per iteration. By contrast, the complexity of the UE localization method in [26] is $\mathcal{O}(\max(K_{\text{I}}N^{3},K_{\text{I}}N_{\text{R}}))$ per iteration. Nevertheless, by adopting an AO-based approach to jointly estimate the UE position, CFO, and PN, our algorithm can achieve up to a two-order-of-magnitude improvement in localization accuracy under high SNR conditions, compared to the method developed in [26], as demonstrated numerically in Section VI-C.
@@
2:  Initialization: $\hat{\phi}$, $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, $k_{\color[rgb]{0,0,0}\text{p}}$, $\epsilon_{\color[rgb]{0,0,0}\text{O}}$, $\epsilon_{\color[rgb]{0,0,0}\text{I}}$, $N_{\color[rgb]{0,0,0}\text{O}}$, $N_{\color[rgb]{0,0,0}\text{I}}$;
@@
5:     Given fixed $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, update $\hat{\bm{\eta}}$ with $\hat{\phi}$ using (15);
@@
7:     Given fixed $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, update $\hat{\phi}$ with $\hat{\bm{\theta}}$ using (22);
@@
9:        $\hat{d}_{\color[rgb]{0,0,0}\text{R}}\leftarrow\hat{d}_{\color[rgb]{0,0,0}\text{R}}-\frac{\partial\mathcal{L}_{\xi}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}k_{\color[rgb]{0,0,0}\text{p}}$;
10:        $\hat{\phi}_{\color[rgb]{0,0,0}\text{az}}\leftarrow\hat{\phi}_{\color[rgb]{0,0,0}\text{az}}-\frac{\partial\mathcal{L}_{\xi}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}k_{\color[rgb]{0,0,0}\text{p}}$;
11:        $\hat{\phi}_{\color[rgb]{0,0,0}\text{el}}\leftarrow\hat{\phi}_{\color[rgb]{0,0,0}\text{el}}-\frac{\partial\mathcal{L}_{\xi}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}k_{\color[rgb]{0,0,0}\text{p}}$;
12:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{\color[rgb]{0,0,0}\text{I}}$ after the $p$-th iteration or $p\textgreater N_{\color[rgb]{0,0,0}\text{I}}$.
13:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{\color[rgb]{0,0,0}\text{O}}$ after the $q$-th iteration or $q\textgreater N_{\color[rgb]{0,0,0}\text{O}}$.
@@
15:  Output: $\hat{\phi}$, $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$.